Solve
(by factoring) x2nd power - 24 = -2x

Answer by stanbon(75887)

Solve
(by factoring) x2nd power - 24 = -2x
---------------------
x^2 + 2x - 24 = 0
(x+6)(x-4) = 0
x = -6 or x = 4
=====================
Cheers,
Stan H.